# Dissipation-Induced Steady States in Topological Superconductors: Mechanisms and Design Principles

### v1

@@ 4.1 The number of ZKM: algebraic approach
In the case $N_{M}>N_{B}$, the number of ZKM is bounded from below as
@@
The equality corresponds to the condition $\operatorname{rk}\mathcal{B}=2N_{B}$ and the following behavior of the characteristic polynomial $\mathcal{P}(\beta\to 0)\sim\beta^{2(N_{M}-N_{B})}$. In this case, we term such ZKM as robust zeroes. In the opposite case $N_{M}\leqslant N_{B}$, the robust ZKM are absent.
@@
However, one can engineer additional $r$ zeroes, $N_{0}=2(N_{M}-N_{B})+r$, due to reduction of the rank of the matrix $\mathcal{B}$: $\operatorname{rk}\mathcal{B}=2\min(N_{B},N_{M})-r$. We will term such additional zeroes as weak zeroes. The straight forward way to maximize the number of ZKM is to engineer the dissipative fields that do not belong to the zero mode subspace of eigen vectors of isolated system. In this way the matrix $\mathcal{B}$ is identical zero such that the dissipation does not reduce the number of zero modes $N_{0}=2N_{M}$ (it can be considered as generalization of the construction proposed in Refs. [68, 69, 70] for the Liouvillians without unitary evolution).
@@ 4.2 The number of ZKM: operator approach and symmetries
In exceptional cases, $N_{0}=0$ ( $\operatorname{rk}\mathcal{B}=2N_{M}$ ) and $N_{0}=1$ ( $\operatorname{rk}\mathcal{B}=2N_{M}-1$ ) the unitary operators in the form of Eq. (59) do not exist. In the case $1<N_{0}<2N_{M}$ the operators $\hat{S}(\,\underline{\theta}\,)$ do exist and are parametrized by $N_{0}(N_{0}-1)/2$ real parameters. In the absence of dissipation, the dimensionality of the parameter space becomes $2N_{M}(2N_{M}-1)/2$ as expected.
@@
$\hat{\omega}_{l}=\underline{y}_{l}\cdot\hat{\underline{b}}\,,\qquad\{\hat{\omega}_{l},\leavevmode\nobreak\ \hat{\omega}_{m}\}=2\delta_{lm},\qquad l,m=1,\ldots,N_{0}$ (61)
@@
${{\Theta}^{(ev)}_{c}}=\bigoplus_{j=1}^{N_{0}/2}\begin{pmatrix}0&i\kappa_{j}\\ -i\kappa_{j}&0\end{pmatrix},\leavevmode\nobreak\ N_{0}\in 2\mathbb{Z}_{+};\quad{{\Theta}^{(od)}_{c}}=\{0\}\!\!\!\!\bigoplus_{j=1}^{(N_{0}-1)/2}\begin{pmatrix}0&i\kappa_{j}\\ -i\kappa_{j}&0\end{pmatrix},\leavevmode\nobreak\ N_{0}\in 2\mathbb{Z}_{+}+1.$ (63)
@@
4.3 The number of ZKM: the structure of eigen vectors
$\ell^{T}\mathcal{B}^{T}\underline{y}_{a}=0,$ (67)
that is trivially satisfied due to $\underline{y}_{a}\in\ker\mathcal{B}^{T}$. Since in Eq. (66) we construct eigen vectors corresponding to all existing ZKM, we can state that there is no other zero eigen vectors of $X$.
The result (66) allows us to put the following physical picture of the effect of dissipation on the MM. Below we assume that $N_{B}<N_{M}$. The first effect is destruction of the $2N_{M}-N_{0}=\operatorname{rk}\mathcal{B}$ number of MM: the corresponding eigen values $\beta_{a}$ becomes nonzero. For the remaining $N_{0}$ ZKM with $\beta_{a}=0$, their eigen vectors become linear combination of the eigen vectors of MM in the isolated system. This fact can be naturally interpreted as hybridization between MM states due to interaction with the bath.
We note the subtlety of the transformation of the eigenvectors of $X$ in the limit of vanishing dissipation, see Sec. 3.5. In the presence of dissipation the eigenvectors of the matrix $X$ for the ZKM not only belongs to the kernel of $A$ but also are real. In the isolated case the eigenvectors of the matrix $A$ for MM are conveniently chosen as complex vectors, $\underline{\chi}_{2a-1}\pm\,i\underline{\chi}_{2a}$, to be consistent with the structure of eigenvectors for states with nonzero energies $\varepsilon_{b}>0$. We note that for the MM in the isolated system we could choose a real eigenvectors of the matrix $A$. Therefore, we will assume the transformation of $U\to\mathcal{W}R$ in the limit of vanishing dissipation (having in mind subtlety discussed above).
@@ 4.4 Recipes to generate weak ZKM
The results above provide the exact expression for the number of robust ZKM: $\max\{2(N_{M}-N_{B}),0\}$. Analysis of the expression (52)-(53) allows us to give several practical recipes for inducing weak ZKM. As we discussed above, if one knows the eigen functions of the zero modes in the isolated system then it is possible to construct dissipative field not belonging to the zero mode subspace such that the total number of ZKM coincides with that in the absence of dissipation. If only some of the dissipative fields are orthogonal to the zero mode eigen functions then we are able to have weak ZKM those number lies between $2N_{M}$ and $2(N_{M}-N_{B})$ (in the case $N_{M}>N_{B}$). We list below several ways to achieve weak ZKM.
@@
    Some dissipative field $l_{v}^{r}\notin\ker A$ (or $l_{v}^{i}\notin\ker A$), i.e. it belongs to linear hull of wave functions $\{\underline{\chi}_{m^{\prime}}\}$ corresponding to the nonzero eigen values of $A$, $\varepsilon_{m^{\prime}}>0$. Since the wave functions are orthogonal, $\underline{\chi}_{n}\cdot\underline{\chi}_{m}=\delta_{n,m}$, the corresponding column of the matrix $\mathcal{B}$ vanishes. Each such dissipative field produces one weak ZKM. Such mechanism of generation of weak ZKM generalizes ideas proposed in Ref. [68, 69, 70] to the case of the presence of unitary evolution in Lindbladian.
@@ 5.1 Formulation of model in the BDI class
We emphasize that exactly $2N_{M}$ Majorana operators ($w_{1}$, $w_{3}$$\ldots,$ $w_{2N_{M}-1}$ on the left part of the chain and $w_{2N-2N_{M}},\dots,w_{2N-2},w_{2N}$ on the right part of the chain) are not involved in the Hamiltonian (74), see Fig.  1b. Therefore, such Majorana operators do commute with the Hamiltonian (74) and realize $2N_{M}$ MM. We mention that a generic Hamiltonian (71) can be adiabatically transformed to the Hamiltonian (74) without closing the bulk gap. We note that the Hamiltonian (74) corresponds to the matrix $A$ in the canonical form, see Eq. (6). Therefore, its eigen vectors are localized at a single site: $\chi_{a,j}=\delta_{ja}$.
@@ 5.3 Results of numerical calculations for the case $N_{M}=N_{B}=1$
The third line of Fig. 2 corresponds to the mechanism of generation of weak ZKM, in which the dissipative fields are orthogonal to one of the Majorana modes, e.g., $\underline{l}^{r,i}\cdot\underline{\chi}_{1}=0$ (see item (b) in Sec. 4.4). In Fig. 2g the dissipative fields were chosen as $\underline{l}^{r,i}=\sum_{s=2}^{6}\,c_{s}^{r,i}\,\underline{\chi}_{s}$ with random coefficients $c^{r,i}\in[0,\leavevmode\nobreak\ 1]$. As can be seen from Fig. 2g and 2h, the spatial structure of dissipative fields in this case may be quite nontrivial, however, for the ZKM eigen vector is $\underline{U}_{1}=\underline{\chi}_{1}$, since the dissipative field does not affect the MM with the wave function $\underline{\chi}_{1}$. An additional feature of the considered situation is the orthogonality of the dissipative fields to the subspace of the other gapped modes with a linear hull $\operatorname{span}\{\,\underline{\chi}_{7},\,\ldots,\,\underline{\chi}_{2N}\}$. Therefore, the eigen values of $X$ includes non-dissipating excitations with $\operatorname{Re}\beta=0$, which correspond to oscillating modes of the isolated system, cf. Fig. 2c and Fig. 2i. This situation corresponds to the fact that the Lindbladian spectrum contains a large number of zero modes, $\lambda_{\underline{\eta}}=0$, see Eq. (28), and NESS is multiply degenerate.

### v2

@@ 4.1 Main result: the number and structure of ZKM
Here, the hybridization matrix $\mathcal{B}$ is constructed from the wave functions of isolated zero modes $\underline{\chi}_{b}$, defined by Eqs. (6)-(7) with $\varepsilon_{b}=0$, and the dissipation fields $\underline{l}_{v}=\underline{l}_{v}^{r}+i\underline{l}_{v}^{i}$ as follows:
$\mathcal{B}=\begin{pmatrix}\mathcal{C}_{1}^{(1)}&\ldots&\mathcal{C}_{1}^{(N_{B})}\\ \vdots&\ddots&\vdots\\ \mathcal{C}_{N_{M}}^{(1)}&\ldots&\mathcal{C}_{N_{M}}^{N_{B}}\end{pmatrix}=\tilde{\mathcal{W}}^{T}\cdot\ell\,\in\,\mathbb{R}^{2N_{M}\times 2N_{B}},\quad\mathcal{C}_{b}^{(v)}=\begin{pmatrix}\underline{\chi}_{2b-1}\cdot\underline{l}^{r}_{v}&\underline{\chi}_{2b-1}\cdot\underline{l}^{i}_{v}\\ \underline{\chi}_{2b}\cdot\underline{l}^{r}_{v}&\underline{\chi}_{2b}\cdot\underline{l}^{i}_{v}\end{pmatrix}$. (44)
The ZKMs themselves are linear combinations of the initially isolated zero modes,
$\underline{U}_{a}=\sum_{b=1}^{2N_{M}}y_{a,b}\,\underline{\chi}_{b}=\tilde{\mathcal{W}}\,\underline{y}_{a},\qquad X\,\underline{U}_{a}=0,\qquad\underline{\chi}_{b}\in\ker A,\qquad\underline{y}_{a}\in\ker\mathcal{B}^{T}.$ (45)
The proof of this theorem is provided in Appendix 7.2.
In Eq.(44) $\tilde{\mathcal{W}}=[\underline{\chi}_{1},\underline{\chi}_{2},\dots,\underline{\chi}_{2N_{M}-1},\underline{\chi}_{2N_{M}}]\in\mathbb{R}^{2N\times 2N_{M}}$ is nothing but the reduced $\mathcal{W}$, cf. Eq. (7), composed out of $2N_{M}$ first columns, corresponding to the zero modes. We note that the matrix $\mathcal{B}$ has the form of cross Gram matrix built out of two sets of vectors $\underline{\chi}_{b}$, $b=1,\dots,2N_{M}$, and $l_{v}^{r,i}$, $v=1,\dots,N_{B}$. The matrix $\mathcal{B}$ has a transparent meaning. The vectors $\underline{\chi}_{a}$, $a=1,\dots,2N$, form a basis in $\mathbb{R}^{2N}$, and the eigen vectors $\chi_{b}$, $b=1,\dots,2N_{M}$, correspond to the zero mode subspace in $\mathbb{R}^{2N}$. The elements of the matrix $\mathcal{B}$ encodes the part of the dissipative field belonging to the zero mode subspace: $\ell=\mathcal{B}^{T}\tilde{\mathcal{W}}$$\ell_{\perp}$.
From (43), if $N_{M}>N_{B}$, then the number of ZKMs is bounded from below as
@@
In this case, we term such ZKM as robust zeroes. In the opposite case $N_{M}\leqslant N_{B}$, the robust ZKM are absent. However, one can engineer additional $r$ zeroes, $N_{0}=2(N_{M}-N_{B})+r$, due to reduction of the rank of the matrix $\mathcal{B}$: $\operatorname{rk}\mathcal{B}=2\min(N_{B},N_{M})-r$. We will term such additional zeroes as weak zeroes, see Sec. 4.5.
The results of Theorem 1 allows us to put the following physical picture of the effect of dissipation on the MM. Below we assume that $N_{B}<N_{M}$. The first effect is destruction of the $2N_{M}-N_{0}=\operatorname{rk}\mathcal{B}$ number of MM: the corresponding eigen values $\beta_{a}$ becomes nonzero. For the remaining $N_{0}$ ZKM with $\beta_{a}=0$, their eigen vectors become linear combination of the eigen vectors of MM in the isolated system. This fact can be naturally interpreted as hybridization between MM states due to interaction with the bath.
@@
The straight forward way to maximize the number of ZKM is to engineer the dissipative fields that do not belong to the zero mode subspace of eigen vectors of isolated system. In this way the matrix $\mathcal{B}$ is identical zero such that the dissipation does not reduce the number of zero modes $N_{0}=2N_{M}$.
We also note another limiting case studied in Refs. [69, 70, 71] where unitary dynamics is absent ($A=0$) in the GKSL equation. Here, the kernel of $A$ has dimension $\dim\ker A=2N$ and can be spanned as $\ker A=\operatorname{span}\{\underline{\chi}_{1},\ldots,\underline{\chi}_{2N}\}$ by Majorana wave functions $\underline{\chi}_{l}$ of an arbitrary single-particle Hamiltonian $A$. Choosing the latter as a topological superconductor Hamiltonian with one pair of MMs $\chi_{1,2}$, and taking the N1N-1 dissipator fields $\underline{l}^{r,i}_{v}$ to be spanned by the gapped modes $\underline{\chi}_{3},\ldots,\underline{\chi}_{2N}$, Theorem 1 enables the realization of two ZKMs with properties analogous to $\chi_{1,2}$, see also Sec. 4.5.
Finally, we remark that Theorem 1 only provides information about the number and structure of the ZKMs, without addressing their topological nature – i.e., whether these modes can be associated with a nontrivial topological index of the corresponding NESS. Nonetheless, partial insights into this matter can already be found in the literature. In particular, the situation was considered in Refs. [69, 70, 71] where, in the absence of unitary evolution, dissipative fields induce a regime corresponding to the symmetric point of the Kitaev chain (see Sec. 5.1). It was shown that the density matrix describing such a NESS can be pure and characterized by a nontrivial topological index. Accordingly, many nontrivial properties of MMs in this construction translate into properties of the ZKMs. In the case of a topological superconductor with spatial dimension $d\geq 2$, addressing this question encounters a no-go theorem, which states the impossibility of realizing a pure NESS with nontrivial topological properties [24].
4.2 Robustness of the ZKM against to interactions and disorder
While Theorem 1 assumes gapless modes in the isolated system ($\varepsilon_{1}=\ldots=\varepsilon_{N_{M}}=0$), a more realistic scenario involves near-zero subgap states with $|\varepsilon_{b}|<\delta\ll\Delta$ ($b=1,\ldots,N_{M}$), where $\Delta$ is the spectral gap. A mathematically rigorous analysis of this situation for an arbitrary number $N_{B}$ of dissipators lies beyond the scope of the present study. However, for the specific case of $N_{B}=1$, the analysis in Appendix 7.3 reveals that the $N_{0}=\max(2N_{M}-2,0)$ robust ZKMs acquire energies $|\beta|\sim\delta$ in the limit $\delta/\Delta\to 0$. This finding demonstrates that the mathematically rigorous assumptions of Theorem 1 remain applicable to physical systems with $\delta/\Delta\ll 1$, provided that terms of order $\sim O(\delta/\Delta)$ are negligible.
This establishes a direct connection between the topological robustness of the MMs and the stability of the corresponding ZKMs in topological superconductors against disorder and interactions. Since Majorana zero modes, which underlie the formation of the near-zero subspace, are robust to moderate disorder and weak electron-electron interactions (see, for example, Refs.[94, 95, 96, 97, 98] for discussions on Majorana wires), the energies of the ZKMs should therefore remain unaffected as well, experiencing only minor shifts in the presence of these factors.
Thus, we conclude that the predictions of Theorem 1 are not merely mathematical idealizations but are expected to hold relevance for realistic topological superconductors subject to disorder and interactions. The key requirement is that the system remains in a topological phase where the subgap modes are only weakly split from zero energy, allowing one to neglect corrections of order $\sim O(\delta/\Delta)$. This provides a robust foundation for anticipating the observable signatures of ZKMs in experimental settings [107].
4.3 The number of ZKM in the presence of non-Markovian reservoirs
Among the dissipative couplings to the considered system, it is worthwhile to single out a practically important case of tunnel interaction with Fermi reservoirs. In general, such a situation cannot be reduced to the GKSL equation. It becomes possible within a number of self-consistent approximations (see Appendix 7.1); however, the analysis of the system’s spectral properties is not straightforward within this framework. Significantly, in the quantum-field approach one can still analyze the ZKM emergence problem even if the memory effects (non-Markovianity) of the reservoirs are involved [107]. Here, in the wide-band limit the reservoirs’ fields included in the characteristic polynomial of $X$ (93) are $\beta$-independent allowing to utilize the Theorem 1 (note the $\beta$-dependence persists in the Keldysh sector, i.e. in the reservoirs’ distribution functions). However, the structure of the these fields differs from Eq. (4). In particular, the corresponding jump operators contain only non-zero $\underline{\mu}_{v}=2\underline{\gamma}_{v}$, while $\underline{\nu}_{v}=0$. Here the vector $\underline{\gamma}_{v}\in\mathbb{C}^{N}$ involves the tunneling amplitudes between the fermionic state at site $l$ and fermionic states in the reservoir $v$. Then, in the presence of $2N_{R}$ reservoirs’ fields among the total $2N_{B}$ dissipative fields, the matrix $\ell$ becomes
${\ell}=[\,\underline{l}_{1}^{r}\,,\,\underline{l}_{1}^{i}\,,\,\ldots\,,\,\underline{l}^{r}_{N_{B}-N_{R}}\,,\,\underline{l}^{i}_{N_{B}-N_{R}},\,\,\underline{t}_{1}^{r}\,,\,\underline{t}_{1}^{i}\,,\,\ldots\,,\,\underline{t}_{N_{R}}^{r}\,,\,\underline{t}_{N_{R}}^{i}\,]\in\mathbb{R}^{2N\times 2N_{B}},$ (47)
where $t_{v,2l-1}=\gamma_{v,l}$ and $t_{v,2l}=i\gamma_{v,l}$ such that
$\begin{split}\underline{t}_{v}^{r}&=(\operatorname{Re}\gamma_{v,1},\,-\operatorname{Im}\gamma_{v,1},\dots,\operatorname{Re}\gamma_{v,N},\,-\operatorname{Im}\gamma_{v,N})^{T},\\ \underline{t}_{v}^{i}&=(\operatorname{Im}\gamma_{v,1},\,\quad\operatorname{Re}\gamma_{v,1},\dots,\operatorname{Im}\gamma_{v,N},\quad\operatorname{Re}\gamma_{v,N})^{T}.\end{split}$ (48)
There is a subtle difference between the dissipative fields $l_{v}^{r,i}$ and the reservoir fields $t_{v}^{r,i}$. Generically, the former are completely independent while the latter are related via the matrix multiplication:
$\underline{t}_{v}^{r}=\bigoplus_{l=1}^{N}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\underline{t}_{v}^{i}\,.$ (49)
Additionally, we mention that vectors $\underline{t}_{v}^{r}$ and $\underline{t}_{v}^{i}$ are orthogonal to each other, $\underline{t}_{v}^{r}\cdot\underline{t}_{v}^{i}=0$, i.e. linear independent. Therefore, such relations between vectors $\underline{t}_{v}^{r}$ and $\underline{t}_{v}^{i}$ do not reduce the rank of the matrix $\mathcal{B}$. Thus the number of robust ZKM is still given by lower bound in Eq. (46).
@@ 4.4 The number of ZKM: operator approach and symmetries
In exceptional cases, $N_{0}=0$ ( $\operatorname{rk}\mathcal{B}=2N_{M}$ ) and $N_{0}=1$ ( $\operatorname{rk}\mathcal{B}=2N_{M}-1$ ) the unitary operators in the form of Eq. (53) do not exist. In the case $1<N_{0}<2N_{M}$ the operators $\hat{S}(\,\underline{\theta}\,)$ do exist and are parametrized by $N_{0}(N_{0}-1)/2$ real parameters. In the absence of dissipation, the dimensionality of the parameter space becomes $2N_{M}(2N_{M}-1)/2$ as expected. Note that in the case of $N_{0}=0$, the non-equilibrium steady state is unique, and the symmetry operators of the form (53) act trivially, i.e., $\hat{S}(\theta)\sim\hat{I}$. Conversely, it is known from the theorems of Refs. [41, 42, 43, 44, 45, 46, 47, 48] that if the relations $\hat{H},\,\hat{S}(\,\underline{\theta}\,)\,]=0$, $\hat{L}_{v},\,\hat{S}(\,\underline{\theta}\,)\,]=0\,$, and $\hat{L}^{+}_{v},\,\hat{S}(\,\underline{\theta}\,)\,]=0\,$ hold only for operators $\hat{S}(\theta)\sim\hat{I}$ [^6] Such a statement is equivalent to the condition that the commutant of the set of operators $\{\hat{H},\hat{L}_{1},\hat{L}_{1}^{+},\ldots,\hat{L}_{N_{B}},\hat{L}_{N_{B}}^{+}\}$ is proportional to the identity operator. The latter is also equivalent to the statement that the set $\{\hat{H},\hat{L}_{1},\hat{L}_{1}^{+},\ldots,\hat{L}_{N_{B}},\hat{L}_{N_{B}}^{+}\}$ generates all the operators under multiplication, addition, and scalar multiplication, see Refs.[45, 46, 48] , then the NESS is unique.
@@
$\hat{\omega}_{l}=\underline{y}_{l}\cdot\hat{\underline{b}}\,,\qquad\{\hat{\omega}_{l},\penalty 10000\ \hat{\omega}_{m}\}=2\delta_{lm},\qquad l,m=1,\ldots,N_{0}$ (55)
@@
${{\Theta}^{(ev)}_{c}}=\bigoplus_{j=1}^{N_{0}/2}\begin{pmatrix}0&i\kappa_{j}\\ -i\kappa_{j}&0\end{pmatrix},\penalty 10000\ N_{0}\in 2\mathbb{Z}_{+};\quad{{\Theta}^{(od)}_{c}}=\{0\}\!\!\!\!\bigoplus_{j=1}^{(N_{0}-1)/2}\begin{pmatrix}0&i\kappa_{j}\\ -i\kappa_{j}&0\end{pmatrix},\penalty 10000\ N_{0}\in 2\mathbb{Z}_{+}+1.$ (57)
@@ 4.5 Recipes to generate weak ZKM
The results above provide the exact expression for the number of robust ZKM: $\max\{2(N_{M}-N_{B}),0\}$. Analysis of the results of Theorem 1 allows us to give several practical recipes for inducing weak ZKM. As we discussed above, if one knows the eigen functions of the zero modes in the isolated system then it is possible to construct dissipative field not belonging to the zero mode subspace such that the total number of ZKM coincides with that in the absence of dissipation. If only some of the dissipative fields are orthogonal to the zero mode eigen functions then we are able to have weak ZKM those number lies between $2N_{M}$ and $2(N_{M}-N_{B})$ (in the case $N_{M}>N_{B}$). We list below several ways to achieve weak ZKM.
@@
    Some dissipative field $l_{v}^{r}\notin\ker A$ (or $l_{v}^{i}\notin\ker A$), i.e. it belongs to linear hull of wave functions $\{\underline{\chi}_{m^{\prime}}\}$ corresponding to the nonzero eigen values of $A$, $\varepsilon_{m^{\prime}}>0$. Since the wave functions are orthogonal, $\underline{\chi}_{n}\cdot\underline{\chi}_{m}=\delta_{n,m}$, the corresponding column of the matrix $\mathcal{B}$ vanishes. Each such dissipative field produces one weak ZKM. Such mechanism of generation of weak ZKM generalizes ideas proposed in Ref. [69, 70, 71] to the case of the absence of unitary evolution in Lindbladian.
@@ 5.1 Formulation of model in the BDI class
We emphasize that exactly $2N_{M}$ Majorana operators ($w_{1}$, $w_{3}$$\ldots,$ $w_{2N_{M}-1}$ on the left part of the chain and $w_{{\color[rgb]{1,0,0}2N-2N_{M}}},\dots,w_{2N-2},w_{2N}$ on the right part of the chain) are not involved in the Hamiltonian (65), see Fig.  1b. Therefore, such Majorana operators do commute with the Hamiltonian (65) and realize $2N_{M}$ MM. We mention that a generic Hamiltonian (62) can be adiabatically transformed to the Hamiltonian (65) without closing the bulk gap. We note that the Hamiltonian (65) corresponds to the matrix $A$ in the canonical form, see Eq. (6). Therefore, its eigen vectors are localized at a single site: $\chi_{a,j}=\delta_{ja}$.
@@ 5.3 Results of numerical calculations for the case $N_{M}=N_{B}=1$
The third line of Fig. 2 corresponds to the mechanism of generation of weak ZKM, in which the dissipative fields are orthogonal to one of the Majorana modes, e.g., $\underline{l}^{r,i}\cdot\underline{\chi}_{1}=0$ (see item (b) in Sec. 4.5). In Fig. 2g the dissipative fields were chosen as $\underline{l}^{r,i}=\sum_{s=2}^{6}\,c_{s}^{r,i}\,\underline{\chi}_{s}$ with random coefficients $c^{r,i}\in[0,\penalty 10000\ 1]$. As can be seen from Fig. 2g and 2h, the spatial structure of dissipative fields in this case may be quite nontrivial, however, for the ZKM eigen vector is $\underline{U}_{1}=\underline{\chi}_{1}$, since the dissipative field does not affect the MM with the wave function $\underline{\chi}_{1}$. An additional feature of the considered situation is the orthogonality of the dissipative fields to the subspace of the other gapped modes with a linear hull $\operatorname{span}\{\,\underline{\chi}_{7},\,\ldots,\,\underline{\chi}_{2N}\}$. Therefore, the eigen values of $X$ includes non-dissipating excitations with $\operatorname{Re}\beta=0$, which correspond to oscillating modes of the isolated system, cf. Fig. 2c and Fig. 2i. This situation corresponds to the fact that the Lindbladian spectrum contains a large number of zero modes, $\lambda_{\underline{\eta}}=0$, see Eq. (28), and NESS is multiply degenerate.